A movie stuntman (mass 80.0 kg) stands on a window
Answer #1

Here ,

let the speed at the floor is v

part A)

at the bottom

speed of hero , u = sqrt(2 * g * h)

u = sqrt(2 * 9.8 * 4)

u = 8.85 m/s

Now ,using conservation of momentum

80 * 8.85= (80 + 70) * v

v = 4.722 m/s

the initial speed of entwined across the floor is 4.72 m/s

part B)

Now , acceleration of bodies ,

a= -u * g

a = -0.265 * 9.8

a = - 2.6 m/s^2

Now, using third equation of motion

v^2 - u^2 = 2 * a * d

- 4.722^2 = = -2 * a 2.6 * d

d = 4.29 m

the distance travelled on the floor is 4.29 m